OpenStudy (anonymous):

Tom and Sam sit down to eat a pie. Tom eats 2/3 of the pie, and Sam eats 1/4 of the pie. What fraction of the pie is left?

OpenStudy (anonymous):

2/3=8/12

OpenStudy (anonymous):

1/4=3/12

OpenStudy (anonymous):

8/12 plus 3/12=11/12

OpenStudy (anonymous):